OpenStudy (anonymous):

Related Rates: If oil spilled from a ruptured tanker spreads in a circle at a rate of 6mi^2 /hr, then how fast is the radius of the circle increasing when the area is 9mi^2? I really need help setting this up correctly.

OpenStudy (anonymous):

\[A(r)=\pi r^2\] \[A'(t)=2\pi r r'(t)\]

OpenStudy (anonymous):

you are told that \[A'(t)=6\] and that and you want \[r'(t)\] when \[A=9\] solve for r from the first equation and then plug in to second to find \[r'(t)\]

OpenStudy (anonymous):

oh awesome I had actually chosen the right formulas for once.

OpenStudy (anonymous):

congrats!

OpenStudy (anonymous):

now if I could just stop second guessing myself.

OpenStudy (anonymous):

thanks for the help.

OpenStudy (anonymous):

yw

OpenStudy (amistre64):

and they wonder why no work gets done ...